# Weighted shift operators on spaces of analytic functions over an annulus - I

### v1

@@ 1. Introduction
$z^{\nu}=(-1)^{\nu}\left(\frac{a_{0}\cdots a_{\nu-1}}{b_{0}\cdots b_{\nu-1}}\right)\sum_{j=-\infty}^{\infty}\lambda_{j}f_{j}-(-1)^{\nu}\left(\frac{a_{1}\cdots a_{\nu-1}}{b_{0}\cdots b_{\nu-1}}f_{0}-\cdots+(-1)^{\nu-1}\frac{1}{b_{\nu-1}}f_{\nu-1}\right).$
Finally, we get
@@
$z^{\nu}$ $=$ $(-1)^{\nu}\left(\frac{1-\lambda_{0}a_{0}}{b_{-1}}\right)\left(\frac{a_{0}\cdots a_{\nu-1}}{b_{0}\cdots b_{\nu-1}}\right)\left(f_{-1}+\sum_{j=2}^{\infty}(-1)^{j-1}\frac{a_{-1}\cdots a_{-j+1}}{b_{-2}\cdots b_{-j}}f_{-j}\right)$
$+(-1)^{\nu}\left(\lambda_{0}a_{0}-1\right)\left(\frac{a_{0}\cdots a_{\nu-1}}{b_{0}\cdots b_{\nu-1}}\right)\left(\frac{1}{a_{0}}f_{0}-\frac{b_{0}}{a_{0}a_{1}}f_{1}+\cdots+(-1)^{\nu-1}\frac{b_{0}\cdots b_{\nu-2}}{a_{0}\cdots a_{\nu-1}}f_{\nu-1}\right)$
$+\frac{\lambda_{0}a_{0}}{a_{\nu}}\left(f_{\nu}+\sum_{j=1}^{\infty}(-1)^{j}\frac{b_{\nu}b_{\nu+1}\cdots b_{\nu+j-1}}{a_{\nu+1}a_{\nu+2}\cdots a_{\nu+j}}f_{\nu+j}\right).$
Now, since $\lim\sup_{n\rightarrow+\infty}|b_{n}/a_{n+1}|<1,$ we can find $r<1$ and an integer $N\geq 1,$ such that $\big{|}\frac{b_{n}}{a_{n+1}}\big{|}<r,$ for all $n\geq N.$ Thus, there exists a constant $M_{1}>0$ such that
$\left\lVert z^{j}\right\rVert_{\ell^{p}_{{a},{b}}(\mathtt{A})}\leq M_{1}\left\lvert\frac{a_{0}a_{1}\cdots a_{j-1}}{b_{0}b_{1}\cdots b_{j-1}}\right\rvert,$
for all $j\geq N$. This proves the first part.
For $\nu\leq-1,$ we write
$b_{-1}=f_{-1}-a_{-1}z^{-1},$
and proceed as above using $\lim\sup_{n\rightarrow+\infty}|a_{-n+1}/b_{-n}|<1.$ Then, we can obtain that
$\left\{\frac{a_{-1}a_{-2}\cdots a_{-\nu}}{b_{-1}b_{-2}\cdots b_{-\nu}}z^{-\nu}:\nu\geq 1\right\}$
is norm-bounded in $\ell^{p}_{a,b}(\mathtt{A}),$ and hence the second estimate in the proposition holds. ∎
@@ 3. Bilateral weighted shifts $F_{w}$ on $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$
In this section we introduce the bilateral forward shift operator $F_{w}$ on the spaces $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$. Necessary and sufficient conditions for $F_{w}$ to become a bounded operator are obtained in terms of $\{a_{n}\}$, $\{b_{n}\}$ and the weight sequence $\{w_{n}\}$. It is shown that operator $F_{w}$, under suitable assumptions, is similar to a compact perturbation of a classical bilateral weighted shift on $\ell^{p}(\mathbb{Z})$. Hypercyclicity, supercyclicity, and chaos for both $F_{w}$ and $F_{w}^{*}$ are studied. Illustrations using examples are provided.
Let $w:=\{w_{n}\}$ be a complex sequence. The operator $F_{w}$ is given by the action $F_{w}(z^{n})=w_{n}z^{n+1}$ for all $n\in\mathbb{Z}$, that is, $F_{w}(\sum_{n=-\infty}^{\infty}\lambda_{n}z^{n})=\sum_{n=-\infty}^{\infty}\lambda_{n}w_{n}z^{n+1}$. We first isolate certain properties of $F_{w}$ which are of general interest. Among them is an important property of weighted forward shifts, which loosely says that the adjoint $F_{w}^{*}$ whenever $F_{w}$ is bounded, acts like a “weighted backward shift” with respect to a total set in the dual of a Banach space of analytic functions. The unweighted case of $F$ defined on reproducing kernel space over the unit disc was discussed in [33].
@@
Note that $k_{n}\in\mathcal{E}^{*}$ for all $n$. For a function $f(z)=\sum_{k=-\infty}^{\infty}\lambda_{k}z^{k}$ in $\mathcal{E}$, we have
@@
$=$ $k_{n}\big{(}\sum_{k=-\infty}^{\infty}\lambda_{k}w_{k}z^{k+1}\big{)}$
$=$ $\lambda_{n-1}w_{n-1}.$
On the other hand, $k_{n-1}(f)=k_{n-1}\left(\sum_{k=-\infty}^{\infty}\lambda_{k}z^{k}\right)=\lambda_{n-1}.$ Hence, $F_{w}^{*}(k_{n})=w_{n-1}k_{n-1},$ for all $n\in\mathbb{Z}.$
@@
Let $F_{w}$ be a bounded forward weighted shift on a Banach space $\mathcal{E}$ consisting of analytic functions on the unit disc $\mathbb{D}$. Then
@@
In terms of $k_{n}$, a general necessary condition for $F_{w}$ to be bounded can be obtained:
Proposition 3.3.
Proof.
3.1. Matrix representation of $F_{w}$, boundedness and a perturbation result
To get the boundedness, we compute the matrix representations of the operator $F_{w}$ acting on $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$ with respect to the ordered Schauder basis $\{f_{n}\}_{n\in\mathbb{Z}}$. Recall that normalized Schauder basis $f_{n}(z)=(a_{n}+b_{n}z)z^{n},n\in\mathbb{Z}$ in $\ell^{p}_{a,b}(\mathtt{A})$ is assumed to be equivalent to the standard basis in $\ell^{p}(\mathbb{Z})$. Then, the boundedness of $F_{w}$ is equivalent to that of its matrix operator acting on $\ell^{p}(\mathbb{Z})$.
Note that
@@ 3.1. Matrix representation of $F_{w}$, boundedness and a perturbation result
by putting the value of $z^{n+1}$. Setting
$c_{n}:=w_{n+1}\frac{b_{n}}{a_{n+2}}-w_{n}\frac{a_{n}}{a_{n+1}}\frac{b_{n+1}}{a_{n+2}},$
we immediately get for $n\geq-1,$
@@
$+(-1)^{n+2}\alpha c_{n}\left(\frac{a_{0}\cdots a_{n+2}}{b_{-1}\cdots b_{n+1}}\right)\left(f_{-1}+\sum_{j=2}^{\infty}(-1)^{j-1}\frac{a_{-1}\cdots a_{-j+1}}{b_{-2}\cdots b_{-j}}f_{-j}\right)$
$+(-1)^{n+3}\alpha c_{n}\left(\frac{a_{1}\cdots a_{n+2}}{b_{0}\cdots b_{n+1}}f_{0}-\frac{a_{2}\cdots a_{n+2}}{b_{1}\cdots b_{n+1}}f_{1}+\cdots+(-1)^{n}\frac{a_{n+1}a_{n+2}}{b_{n}b_{n+1}}f_{n}\right)$
@@
Again
$F_{w}(f_{n})(z)=w_{n}a_{n}z^{n+1}+w_{n+1}b_{n}z^{n+2}=\frac{w_{n+1}b_{n}}{b_{n+1}}f_{{n+1}}+\left(\frac{w_{n}a_{n}}{b_{n}}-\frac{w_{n+1}a_{n+1}}{b_{n+1}}\right)b_{n}z^{n+1},$
by substituting the expression of $z^{n+2}$. Setting
@@
$+(-1)^{n-1}(1-\alpha)d_{-n}\left(\frac{-b_{-2}\cdots b_{-n}}{a_{-1}\cdots a_{-n+1}}f_{-1}+\cdots+(-1)^{n-2}\frac{b_{-n+1}b_{-n}}{a_{-n+2}a_{-n+1}}f_{-n+2}\right)$
$+(-1)^{n-1}d_{-n}\lambda_{0}\left(\frac{b_{-1}\cdots b_{-n}}{a_{-1}\cdots a_{-n+1}}\right)\left(f_{0}+\sum_{j=1}^{\infty}(-1)^{j}\frac{b_{0}b_{1}\cdots b_{j-1}}{a_{1}a_{2}\cdots a_{j}}f_{j}\right),$
where $\alpha=1-\lambda_{0}a_{0}.$ Hence, the (bilateral) matrix $[F_{w}]$ of the operator $F_{w}$ with respect to the ordered normalized Schauder basis $\{f_{n}\}_{n\in\mathbb{Z}}$ is given by
@@
$\text{diag}~{}(\cdots,\alpha d_{-3},\alpha d_{-2},\alpha\frac{c_{-1}a_{0}a_{1}}{b_{-1}b_{0}},\bm{\alpha\frac{c_{0}a_{1}a_{2}}{b_{0}b_{1}}},\alpha\frac{c_{1}a_{2}a_{3}}{b_{1}b_{2}},\alpha\frac{c_{2}a_{3}a_{4}}{b_{2}b_{3}},\alpha\frac{c_{3}a_{4}a_{5}}{b_{3}b_{4}},\cdots)$
on $\ell^{p}(\mathbb{Z}).$ Also the matrix $[E_{n}]$ is obtained by deleting all the entries of $[F_{w}]$, except those at the $n$-th subdiagonal, where $n\geq 1$. Observe that $[E_{n}]$ is the matrix of suitable powers of a bilateral weighted forward shift $E_{n}$ for $n\geq 1$. Again, the matrix $[E_{-n}]$ is obtained by deleting all the entries of $[F_{w}]$, except those at the $n$-th superdiagonal, where $n\geq 1$. Observe that $[E_{-n}]$ is the matrix of suitable powers of a bilateral weighted backward shift $E_{-n}$ for $n\geq 1$.
Note that $[E_{1}]$ is the matrix of the standard weighted bilateral forward shift $E_{1}(e_{n})\mapsto\gamma_{n}e_{n+1}$ on $\ell^{p}(\mathbb{Z})$, $n\in\mathbb{Z}$, having weights
@@
By the assumptions, it follows that $\{c_{n}\}$ and $\{d_{n}\}$ is a bounded sequence. Now
$\lVert D\rVert$$\leq$ $\alpha\sup_{n\geq 0}\left\{\left\lvert c_{n}\frac{a_{n+1}a_{n+2}}{b_{n}b_{n+1}}\right\rvert,|d_{-n-2}|\right\},$
$\sum_{i=1}^{\infty}\lVert E_{i}\rVert$$\leq$ $\sup_{n\geq 0}\left\{\left\lvert w_{n}\frac{a_{n}}{a_{n+1}}+\alpha\frac{c_{n}a_{n+2}}{b_{n+1}}\right\rvert,\left\lvert w_{-n-1}\frac{b_{-n-1}}{b_{-n}}+(1-\alpha)\frac{d_{-n-2}b_{-n-2}}{a_{-n-1}}\right\rvert\right\}$
$+(1-\alpha)\sup_{n\geq 0}\left\{\left\lvert c_{n}\right\rvert,\left\lvert\frac{d_{-n-3}b_{-n-2}b_{-n-3}}{a_{-n-1}a_{-n-1}}\right\rvert\right\}$
$+(1-\alpha)\sup_{n\geq 0}\left\{\left\lvert\frac{c_{n}b_{n+2}}{a_{n+3}}\right\rvert,\left\lvert\frac{d_{-n-4}b_{-n-2}b_{-n-3}b_{-n-4}}{a_{-n-1}a_{-n-2}a_{-n-3}}\right\rvert\right\}$
$+(1-\alpha)\sup_{n\geq 0}\left\{\left\lvert\frac{c_{n}b_{n+2}b_{n+3}}{a_{n+3}a_{n+4}}\right\rvert,\left\lvert\frac{d_{-n-5}b_{-n-2}b_{-n-3}b_{-n-4}b_{-n-5}}{a_{-n-1}a_{-n-2}a_{-n-3}a_{-n-4}}\right\rvert\right\}+\cdots,$
@@
$\sum_{i=1}^{\infty}\lVert E_{-i}\rVert$$\leq$ $\alpha\sup_{n\geq 0}\left\{\left\lvert\frac{c_{n}a_{n}a_{n+1}a_{n+2}}{b_{n-1}b_{n}b_{n+1}}\right\rvert,\left\lvert d_{-n-2}\frac{a_{-n-2}}{b_{-n-3}}\right\rvert\right\}$
$+\alpha\sup_{n\geq 0}\left\{\left\lvert\frac{c_{n+1}a_{n}a_{n+1}a_{n+2}a_{n+3}}{b_{n-1}b_{n}b_{n+1}b_{n+2}}\right\rvert,\left\lvert d_{-n-2}\frac{a_{-n-2}a_{-n-3}}{b_{-n-2}b_{-n-3}}\right\rvert\right\}$
$+\alpha\sup_{n\geq 0}\left\{\left\lvert\frac{c_{n+2}a_{n}a_{n+1}a_{n+2}a_{n+3}a_{n+4}}{b_{n-1}b_{n}b_{n+1}b_{n+2}b_{n+3}}\right\rvert,\left\lvert d_{-n-2}\frac{a_{-n-2}a_{-n-3}a_{-n-4}}{b_{-n-3}b_{-n-4}b_{-n-5}}\right\rvert\right\}+\cdots.$
Moreover, our assumptions
$\limsup_{n\rightarrow+\infty}\left\{\left\lvert\frac{b_{n}}{a_{n+1}}\right\rvert,\left\lvert\frac{a_{-n+1}}{b_{-n}}\right\rvert,\left\lvert c_{n}\frac{a_{0}a_{1}\cdots a_{n+2}}{b_{0}b_{1}\cdots b_{n+1}}\right\rvert,\left\lvert d_{-n}\frac{b_{-1}b_{-2}\cdots b_{-n}}{a_{-1}a_{-2}\cdots a_{-n+1}}\right\rvert\right\}<1$
implies that there exists $r,R,s,S<1$ and $n_{0}\in\mathbb{N}$ such that
[…]
for $n\geq n_{0},$ from which it follows that $\sum_{i=1}^{\infty}\lVert E_{-i}\rVert+\lVert D\rVert+\sum_{i=1}^{\infty}\lVert E_{i}\rVert<\infty$ and this completes the proof of the theorem. ∎
@@
The matrix representation of $F_{w}$ hints that the operator $F_{w}$ is far from being similar to a usual weighted shift on $\ell^{p}(\mathbb{Z})$ in general. Indeed, we can produce a weight $w$ and some $a=\{a_{n}\}$ and $b=\{b_{n}\}$ such that the $F_{w}$ acting on $\ell^{p}_{a,b}(\mathtt{A})$ is not similar any forward weighted shift on $\ell^{p}(\mathbb{Z})$, see Section $5$. Below, we provide sufficient conditions so that the operator becomes a compact perturbation of a usual weighted shift.

### v2

@@ 1. Introduction
$\left\lVert z^{\nu}\right\rVert^{p}$ $=$ $\left\lvert(1-\lambda_{0}a_{0})\prod_{k=0}^{\nu-1}\frac{a_{k}}{b_{k}}\right\rvert^{p}\left(\left\lvert\frac{1}{b_{-1}}\right\rvert^{p}+\sum_{j=2}^{\infty}\left\lvert\frac{1}{b_{-1}}\prod_{k=1}^{j-1}\frac{a_{-k}}{b_{-k-1}}\right\rvert^{p}+\left\lvert\frac{1}{a_{0}}\right\rvert^{p}+\sum_{j=1}^{\nu-1}\left\lvert\frac{1}{a_{0}}\prod_{k=1}^{j}\frac{b_{k-1}}{a_{k}}\right\rvert^{p}\right)$
@@
$+\left\lvert\frac{\lambda_{0}a_{0}}{a_{\nu}}\right\rvert^{p}\left(1+\sum_{j=1}^{\infty}\left\lvert\prod_{k=0}^{j-1}\frac{b_{\nu+k}}{a_{\nu+k+1}}\right\rvert^{p}\right).$
We immediately calculate the norm estimates of $z^{\nu}$ using the choices of $N_{1}$ and $N_{2}$. Thus, there exists a constant $M_{1}>0$ such that
$\left\lVert z^{\nu}\right\rVert_{\ell^{p}_{{a},{b}}(\mathtt{A})}\leq M_{1}\left\lvert\frac{a_{0}a_{1}\cdots a_{\nu-1}}{b_{0}b_{1}\cdots b_{\nu-1}}\right\rvert,$
for all $\nu\geq N_{1}$. This is the first estimate in the proposition.
To get the second estimate, we write
$b_{-1}$ $=$ $f_{-1}-a_{-1}z^{-1}$
$=$ $f_{-1}-\frac{a_{-1}}{b_{-2}}f_{-2}+\frac{a_{-1}a_{-2}}{b_{-2}}z^{-2}$
$=$ $f_{-1}+\sum_{j=2}^{\nu}(-1)^{j-1}\frac{a_{-1}a_{-2}\cdots a_{-j+1}}{b_{-2}b_{-3}\cdots b_{-j}}f_{-j}+(-1)^{\nu}\frac{a_{-1}\cdots a_{-\nu}}{b_{-2}\cdots b_{-\nu}}z^{-\nu},$
for a fixed $\nu>1.$ It follows that by using the expansion of the constant $1,$
$z^{-\nu}=(-1)^{\nu}\frac{b_{-1}\cdots b_{-\nu}}{a_{-1}\cdots a_{-\nu}}\left(\sum_{j=-\infty}^{\infty}\lambda_{j}f_{j}\right)-(-1)^{\nu}\left(\frac{b_{-2}\cdots b_{-\nu}}{a_{-1}\cdots a_{-\nu}}f_{-1}-\cdots+(-1)^{\nu-1}\frac{1}{a_{-\nu}}f_{-\nu}\right).$
Thus, we get
$z^{-\nu}$ $=$ $\frac{1-\lambda_{0}a_{0}}{b_{-\nu-1}}\left(f_{-\nu-1}+\sum_{j=0}^{\infty}(-1)^{j+1}\frac{a_{-\nu-1}\cdots a_{-\nu-1-j}}{b_{-\nu-2}\cdots b_{-\nu-2-j}}f_{-\nu-2-j}\right)$
$+(-1)^{\nu}\lambda_{0}a_{0}\frac{b_{-1}\cdots b_{-\nu}}{a_{-1}\cdots a_{-\nu}}\left(\frac{-1}{b_{-1}}f_{-1}+\sum_{j=2}^{\nu}(-1)^{j}\frac{a_{-1}\cdots a_{-j+1}}{b_{-1}\cdots b_{-j}}f_{-j}\right)$
$+(-1)^{\nu}\lambda_{0}\frac{b_{-1}\cdots b_{-\nu}}{a_{-1}\cdots a_{-\nu}}\left(f_{0}+\sum_{j=1}^{\infty}(-1)^{j}\frac{b_{0}b_{1}\cdots b_{j-1}}{a_{1}a_{2}\cdots a_{j}}f_{j}\right),$
which is the expansion of $z^{-\nu}$ in terms of the basis $\{f_{j}\}.$ Computing the norm $\|z^{-\nu}\|$ using the choices of $N_{1}$ and $N_{2},$ the second estimate in the proposition holds. ∎
@@ 3. Bilateral weighted shifts $F_{w}$ on $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$
In this section we introduce the bilateral weighted forward shift operator $F_{w}$ on the spaces $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$. We find boundedness conditions for $F_{w}$. The essential spectrum of $F_{w}$ is obtained via a compact perturbation result. Further, hypercyclicity, supercyclicity, and chaos for both $F_{w}$ and $F_{w}^{*}$ are studied along with examples.
Let $w:=\{w_{n}\}$ be a complex sequence. The operator $F_{w}$ is defined by
$F_{w}(\sum_{n=-\infty}^{\infty}\lambda_{n}z^{n}):=\sum_{n=-\infty}^{\infty}\lambda_{n}w_{n}z^{n+1}.$
So, we have $F_{w}(z^{n})=w_{n}z^{n+1}$ for all $n\in\mathbb{Z}$. We isolate certain general properties of $F_{w}$ which are of independent interest. Among them is an important property of weighted forward shifts, which loosely says that the adjoint $F_{w}^{*}$ whenever $F_{w}$ is bounded, “acts like” a weighted backward shift with respect to a total set, as proved below. The unweighted case of $F$ defined on reproducing kernel space over the unit disc was discussed in [32].
@@
We have already showed that $k_{n}\in\mathcal{E}^{*}$ for all $n$. For a function $f(z)=\sum_{k=-\infty}^{\infty}\lambda_{k}z^{k}$ in $\mathcal{E}$, we have $k_{n-1}(f)=k_{n-1}\left(\sum_{k=-\infty}^{\infty}\lambda_{k}z^{k}\right)=\lambda_{n-1}.$ On the other hand,
@@
$=$ $k_{n}\big{(}\sum_{k=-\infty}^{\infty}\lambda_{k}w_{k}z^{k+1}\big{)}=\lambda_{n-1}w_{n-1}.$
@@
Let $F_{w}$ be a bounded forward weighted shift on a Banach space $\mathcal{E}$ consisting of analytic functions on the unit disc $\mathbb{D}$, having continuous evaluation functionals. Then,
@@
3.1. The matrix of $F_{w}$, boundedness, compact perturbations, and the essential spectrum
To derive some necessary and sufficient conditions for the weighted forward shift $F_{w}$ to be bounded, we compute the matrix representation of $F_{w}$ acting on $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$ with respect to the ordered normalized Schauder basis $f_{n}(z)=(a_{n}+b_{n}z)z^{n},n\in\mathbb{Z}$ in $\ell^{p}_{a,b}(\mathtt{A})$. Then, the boundedness of $F_{w}$ is equivalent to that of its matrix operator acting on $\ell^{p}(\mathbb{Z})$.
We will use the basis expansions of $z^{n}$ given in the proof of Proposition 2.4. To find the matrix of $F_{w}$, we have
@@ 3.1. The matrix of $F_{w}$, boundedness, compact perturbations, and the essential spectrum
by substituting $z^{n+1}=\frac{1}{a_{n+1}}(f_{n+1}-b_{n+1}z^{n+2})$. Set
$c_{n}:=w_{n+1}\frac{b_{n}}{a_{n+2}}-w_{n}\frac{a_{n}}{a_{n+1}}\frac{b_{n+1}}{a_{n+2}}.$
In the following, we now get the $n$-th column of the matrix of $F_{w}$, for $n\geq-1$, by using the expansion of $z^{n+2}$ in terms of the basis elements $f_{n}$’s:
@@
$+(-1)^{n+2}\alpha c_{n}\frac{a_{0}\cdots a_{n+2}}{b_{-1}\cdots b_{n+1}}\left(f_{-1}+\sum_{j=2}^{\infty}(-1)^{j-1}\frac{a_{-1}\cdots a_{-j+1}}{b_{-2}\cdots b_{-j}}f_{-j}\right)$
$+(-1)^{n+3}\alpha c_{n}\left(\sum_{j=1}^{n+1}(-1)^{j-1}\frac{a_{n+2}a_{n+1}\cdots a_{j}}{b_{n+1}b_{n}\cdots b_{j-1}}f_{j-1}\right)$
@@
Similarly, by substituting the expression $z^{n+2}=\frac{1}{b_{n+1}}(f_{n+1}-a_{n+1}z^{n+1})$, we find that
$F_{w}(f_{n})(z)=w_{n}a_{n}z^{n+1}+w_{n+1}b_{n}z^{n+2}=\frac{w_{n+1}b_{n}}{b_{n+1}}f_{{n+1}}+\left(\frac{w_{n}a_{n}}{b_{n}}-\frac{w_{n+1}a_{n+1}}{b_{n+1}}\right)b_{n}z^{n+1}.$
Setting
@@
$+(-1)^{n-1}(1-\alpha)d_{-n}\left(\sum_{j=1}^{n-2}(-1)^{j}\frac{b_{-n}b_{-n+1}\cdots b_{-j-1}}{a_{-n+1}a_{-n+2}\cdots a_{-j}}f_{-j}\right)$
$+(-1)^{n-1}d_{-n}\lambda_{0}\frac{b_{-1}\cdots b_{-n}}{a_{-1}\cdots a_{-n+1}}\left(f_{0}+\sum_{j=1}^{\infty}(-1)^{j}\frac{b_{0}b_{1}\cdots b_{j-1}}{a_{1}a_{2}\cdots a_{j}}f_{j}\right),$
where $n\geq 2$, and $\alpha=1-\lambda_{0}a_{0}.$ Hence, the (bilateral) matrix $[F_{w}]$ of the operator $F_{w}$ with respect to the ordered normalized Schauder basis $\{f_{n}\}_{n\in\mathbb{Z}}$ is given by
@@
$\text{diag}~{}(\cdots,\alpha d_{-3},\alpha d_{-2},-\alpha\frac{c_{-1}a_{0}a_{1}}{b_{-1}b_{0}},\bm{-\alpha\frac{c_{0}a_{1}a_{2}}{b_{0}b_{1}}},-\alpha\frac{c_{1}a_{2}a_{3}}{b_{1}b_{2}},-\alpha\frac{c_{2}a_{3}a_{4}}{b_{2}b_{3}},-\alpha\frac{c_{3}a_{4}a_{5}}{b_{3}b_{4}},\cdots)$
on $\ell^{p}(\mathbb{Z}).$ Also, the matrix $[E_{i}]$ is obtained by deleting all the entries of $[F_{w}]$, except those at the $i$-th subdiagonal, where $i\geq 1$. Observe that $[E_{i}]$ is the matrix of suitable powers of a bilateral weighted forward shift. On the other hand, the matrix $[E_{-i}]$ is obtained by deleting all the entries of $[F_{w}]$, except those at the $i$-th superdiagonal, where $i\geq 1$. Again observe that, $[E_{-i}]$ is the matrix of a suitable power of a bilateral weighted backward shift. Note that $[E_{1}]$ is the matrix of the standard weighted bilateral forward shift $E_{1}(e_{n})\mapsto\gamma_{n}e_{n+1}$ on $\ell^{p}(\mathbb{Z})$, $n\in\mathbb{Z}$, having weights
@@
By the assumptions, it follows that $\{c_{n}\}$ and $\{d_{n}\}$ are bounded sequences. Moreover, we formally have
$\lVert D\rVert\leq\lvert\alpha\rvert\sup_{n\geq 0}\left\{\left\lvert c_{n}\frac{a_{n+1}a_{n+2}}{b_{n}b_{n+1}}\right\rvert,|d_{-n-2}|\right\},$
$\lVert E_{1}\rVert\leq\sup_{n\in\mathbb{Z}}~{}~{}\lvert\gamma_{n}\rvert,~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\lVert E_{2}\rVert\leq|1-\alpha|\sup_{n\geq 0}\left\{\left\lvert c_{n}\right\rvert,\left\lvert\frac{d_{-n-3}b_{-n-2}b_{-n-3}}{a_{-n-1}a_{-n-2}}\right\rvert\right\},$
$\lVert E_{i}\rVert\leq|1-\alpha|\sup_{n\geq 0}\left\{\left\lvert c_{n}\prod_{k=2}^{i-1}\frac{b_{n+k}}{a_{n+k+1}}\right\rvert,\left\lvert d_{-n-i-1}\prod_{k=2}^{i+1}\frac{b_{-n-k}}{a_{-n-k+1}}\right\rvert\right\},~{}~{}~{}~{}\forall i\geq 3,$
@@
$\lVert E_{-i}\rVert\leq|\alpha|\sup_{n\geq 0}\left\{\left\lvert c_{n+i-1}\prod_{k=0}^{i+1}\frac{a_{n+k}}{b_{n+k-1}}\right\rvert,\left\lvert d_{-n-2}\prod_{k=2}^{i+1}\frac{a_{-n-k}}{b_{-n-k-1}}\right\rvert\right\},~{}~{}~{}~{}\forall i\geq 1.$
On the other hand, by the conditions in the statement of the theorem, namely
$\limsup_{n\rightarrow+\infty}\left\{\left\lvert\frac{b_{n}}{a_{n+1}}\right\rvert,\left\lvert\frac{a_{-n+1}}{b_{-n}}\right\rvert,\left\lvert c_{n}a_{0}\prod_{k=1}^{n+2}\frac{a_{k}}{b_{k-1}}\right\rvert,\left\lvert d_{-n}b_{-1}\prod_{k=1}^{n-1}\frac{b_{-k-1}}{a_{-k}}\right\rvert\right\}<1,$ there are $r,R,s,S<1$ and $n_{0}\in\mathbb{N}$ such that
$\left\lvert\frac{b_{n}}{a_{n+1}}\right\rvert<r,\hskip 5.69046pt\left\lvert\frac{a_{-n+1}}{b_{-n}}\right\rvert<R,\hskip 5.69046pt\left\lvert c_{n}a_{0}\prod_{k=1}^{n+2}\frac{a_{k}}{b_{k-1}}\right\rvert<s,\hskip 5.69046pt\text{and}\hskip 5.69046pt\left\lvert d_{-n}b_{-1}\prod_{k=1}^{n-1}\frac{b_{-k-1}}{a_{-k}}\right\rvert<S,$
for $n\geq n_{0},$ from which it follows that
$\lVert E_{i}\rVert\leq|1-\alpha|\max\left\{r^{i-1},SR^{n-n_{0}+2}\right\},~{}~{}\forall i\geq 3~{}~{}\text{and}~{}~{}\lVert E_{-i}\rVert\leq|\alpha|\max\left\{sr^{n-n_{0}},R^{i+1}\right\},~{}~{}\forall i\geq 1.$
Hence, the series $\sum_{i=1}^{\infty}\lVert E_{-i}\rVert$ and $\sum_{i=1}^{\infty}\lVert E_{i}\rVert$ are convergent. Consequently, the matrix $[F_{w}]$ acts as a bounded operator on $\ell^{p}(\mathbb{Z})$. This completes the proof of the theorem. ∎
@@
In terms of $k_{n}$, a general necessary condition for $F_{w}$ to be bounded can be obtained, as follows:
Proposition 3.5.
Proof.
The matrix representation of $F_{w}$ hints that the operator $F_{w}$ is far from being similar to a usual weighted shift on $\ell^{p}(\mathbb{Z})$ in general. Indeed, we can produce a weight $w$, some $a=\{a_{n}\}$ and $b=\{b_{n}\}$ such that the $F_{w}$ acting on $\ell^{p}_{a,b}(\mathtt{A})$ is not similar to any forward weighted shift on $\ell^{p}(\mathbb{Z})$, see Section $5$. Below, we provide sufficient conditions so that the operator becomes a compact perturbation of a usual weighted shift. A consequence is the description of the essential spectrum $\sigma_{e}(F_{w})$ of the weighted shift $F_{w}$ acting on $\ell^{p}_{a,b}(\mathtt{A})$, for $1<p<\infty$. For a bounded operator $T$ on a complex Banach space $X$, the essential spectrum is denoted and defined by
$\sigma_{e}(T):=\{\lambda\in\mathbb{C}:$   $T-\lambda I$ is not Fredholm}}\}},
cf. Shields [36]. See, also, Bayart and Matheron [4], and Douglas [17].
@@
Consequently, $\sigma_{e}(F_{w})=\sigma_{e}(E_{\alpha})$ which is an annulus.